OpenStudy (anonymous):

Could any1 show me the steps to solve: 37x+102.5=-4x

OpenStudy (anonymous):

Simplifying 37x + 102.5 = -4x Reorder the terms: 102.5 + 37x = -4x Solving 102.5 + 37x = -4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '4x' to each side of the equation. 102.5 + 37x + 4x = -4x + 4x Combine like terms: 37x + 4x = 41x 102.5 + 41x = -4x + 4x Combine like terms: -4x + 4x = 0 102.5 + 41x = 0 Add '-102.5' to each side of the equation. 102.5 + -102.5 + 41x = 0 + -102.5 Combine like terms: 102.5 + -102.5 = 0.0 0.0 + 41x = 0 + -102.5 41x = 0 + -102.5 Combine like terms: 0 + -102.5 = -102.5 41x = -102.5 Divide each side by '41'. x = -2.5